# Sensing-Enhanced Channel Estimation for Near-Field XL-MIMO Systems

### v3

@@ II-C Sensing Model
${~{}~{}~{}}\!+\!\sum_{\ell=1}^{L}\!\sum_{m=1}^{N_{\rm UE}}\!\!\alpha_{\ell}{g}\left({\bf r}_{\rm S}^{(\ell)},{\bf r}_{\rm BS}^{(n)}\right){g}\left({\bf r}_{\rm S}^{(\ell)},{\bf r}_{\rm UE}^{(m)}\right){\bf s}_{\rm S}[m],$
where ${\bf s}_{\rm S}\in\mathbb{C}^{N_{\rm UE}}$ denotes the sensing signal. Without loss of generality, we consider the case when ${\bf s}_{\rm S}$ is an all-one vector [26]. As depicted in Fig. 3, the received object waves are then combined with the reference wave as ${\bf e}_{\rm ref}[n]+{\bf e}_{\rm obj}[n]$, where ${\bf e}_{\rm ref}=A[e^{\jmath\xi_{1}},\cdots,e^{\jmath\xi_{N_{\rm BS}}}]^{T}$ is the phase-shifted signal from the local oscillator with $A$ denoting the manipulated amplitude of the reference wave and $\xi_{n}$ denoting the phase shift at the $n$-th antenna element of the BS array. Then, the $n$-th power sensor [31, 32] deployed at the XL-MIMO transceiver measures the power pattern of combined signals as
@@ II-D Pilot Training Model
where ${\bf W}_{\rm RF}^{(t)}\in\mathbb{C}^{N_{\rm BS}\times N^{\rm RF}_{\rm BS}}$ and ${\bf W}_{\rm BB}^{(t)}\in\mathbb{C}^{N^{\rm RF}_{\rm BS}\times N^{\rm S}_{\rm BS}}$ denote the analog and digital combining matrices, and ${\bf F}_{\rm RF}^{(t)}\in\mathbb{C}^{N_{\rm UE}\times N^{\rm RF}_{\rm UE}}$ and ${\bf F}_{\rm BB}^{(t)}\in\mathbb{C}^{N^{\rm RF}_{\rm UE}\times N^{\rm S}_{\rm UE}}$ denote the analog and digital precoding matrices, respectively. Here, $N_{\rm UE}^{\rm RF}$ ($N_{\rm BS}^{\rm RF}$) and $N_{\rm UE}^{\rm S}$ ($N_{\rm BS}^{\rm S}$) denote the numbers of RF chains and data streams at the UE (BS), respectively, ${\bf n}^{(t)}\sim\mathcal{CN}(0,\sigma_{\rm n}^{2}{\bf I})$ is the additive white Gaussian noise (AWGN) vector, and $\sigma_{\rm n}^{2}$ is the power of the AWGN at each receive antenna.
@@ IV-A Time Inversion Near-Field Reconstruction
$+\left.\frac{e^{-2\jmath\xi_{n}+\jmath\kappa\left(\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|+\|{\bf r}-{\bf r}_{\rm BS}^{(n)}\|\right)}}{\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|}\right)$
@@
$\tilde{E}_{\ell}^{\rm d}({\bf r})=\sum_{n=1}^{N_{\rm BS}}\frac{e^{-\jmath\kappa\left(\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|-\|{\bf r}-{\bf r}_{\rm BS}^{(n)}\|\right)}}{\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|}$ (18)
@@
$\tilde{E}_{\ell}^{\rm i}({\bf r})=\sum_{n=1}^{N_{\rm BS}}\frac{e^{-2\jmath\xi_{n}+\jmath\kappa\left(\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|+\|{\bf r}-{\bf r}_{\rm BS}^{(n)}\|\right)}}{\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|}$ (19)
denote the desired term that contributes to the reconstruction and the interference term introduced by the power sampling procedure (7) of the $\ell$-th path, respectively.
@@
$\lim_{\begin{subarray}{c}\kappa\to{\infty}\\ N_{\rm BS}\to{\infty}\end{subarray}}\!\!\left|\tilde{E}_{\ell}^{\rm d}({\bf r})\right|\!=\!\begin{cases}A^{2}N_{\rm BS}^{2}\sum_{n=1}^{N_{\rm BS}}\frac{1}{\left(d^{(\ell,n)}\right)^{2}}\!\!\!\!&\text{if}~{}{\bf r}={\bf r}_{\rm S}^{(\ell)}\!,\\ 0&\text{otherwise}.\end{cases}$ (22)
and the same for the interference term as
@@ IV-B Low-Complexity Time Inversion Algorithm
The locations of the UE and scatterers can be reconstructed exploiting the proposed time inversion method by enumerating possible target locations ${\bf r}$ in the near-field grid. Specifically, during the sensing stage, the BS array generates a uniformly distributed vector $\bm{\xi}=[\xi_{1},\cdots,\xi_{N_{\rm BS}}]$ and performs phase shift $e^{\jmath\xi_{n}}$ at the $n$-th antenna element. The near-field region is split into $G_{x}\times G_{y}$ grids, where $G_{x}$ ($G_{x}>N_{\rm BS}$) and $G_{y}$ denote the numbers of grids on the $x$-axis and the $y$-axis, respectively. The compensated wave at the $i$-th grid on $x$-axis and the $j$-th grid on $y$-axis, i.e., ${\bf r}^{(i,j)}=(x_{i},y_{j})$, is finally manipulated by the time inversion method using (14), and is then stored into the $i$-th row and the $j$-th column of the amplitude matrix $\bm{\Theta}\in\mathbb{R}^{G_{x}\times G_{y}}$, where
@@
which can be reformulated as the linear convolution of vector ¯$\overline{{\bm{\hbar}}}={{\bm{\hbar}}}\odot{\bf e}_{\rm ref}\in\mathbb{C}^{N_{\rm BS}}$ defined in (12) and $\overline{\bf g}_{j}=[e^{\jmath\kappa\sqrt{x_{1}^{2}+y_{i}^{2}}},\cdots,e^{\jmath\kappa\sqrt{x_{G_{x}}^{2}+y_{i}^{2}}}]^{T}\in\mathbb{C}^{G_{x}\times 1}$. Hence, for a fixed $j$-th grid on $y$-axis, the time inversion reconstructed result [^6] Convolution extends the dimension to $N_{\rm BS}+G_{x}-1$, which leads to incomplete reconstruction at the edges, and the valid results are obtained only for indices from $N_{\rm BS}$ to $G_{x}$. To satisfy the need for $G_{x}$ valid points, the $x$-axis grid will be extended to $G_{x}^{\prime}=G_{x}+2N_{\rm BS}$ to compensate the associated edge effects. can be recast as
@@
where $\overline{\bf g}_{j}=[e^{\jmath\kappa\sqrt{x_{1}^{2}+y_{i}^{2}}},\cdots,e^{\jmath\kappa\sqrt{x_{G_{x}^{\prime}}^{2}+y_{i}^{2}}}]^{T}\in\mathbb{C}^{G_{x}^{\prime}\times 1}$ with $G_{x}^{\prime}=G_{x}+2N_{\rm BS}$. Notably, (28) can be further implemented efficiently using FFT as
@@ IV-C Proposed Multipath Localization
$\hat{\alpha}_{\ell}=\overline{\alpha}_{\ell}N_{\rm BS}\sqrt{\sum_{n=1}^{N_{\rm BS}}\frac{1}{\left(d^{(\ell,n)}\right)^{2}}}.$ (30)
@@ VI-C Sensing-Enhanced CE
We firstly investigate the sparsification performance of the proposed dictionary. The channel sparse representations of the DFT dictionary, spherical wave dictionary, and proposed dictionary are compared in Fig. 6(a), where the sparse representation is obtained by $\tilde{\bf h}=\bm{\Psi}^{\dagger}{\bf h}$. As can be observed, the conventional DFT dictionary shows a severe energy leakage effect in the near-field region [13], which can be improved by the spherical wave dictionary sampled in the polar-domain. Meanwhile, the proposed DPSS-based eigen-dictionary entails the sparsest pattern among the three dictionaries. Note that the proposed dictionary is compensated by the matrix ${\bf D}_{{\rm BS}({\rm UE})}$ in (37) and therefore, the sparse representation contains no specific angular information. In addition, different from the DFT and spherical dictionaries, the non-zero support set appears in the first several indices since the SVD always sorts the non-zero singular values first. We further validate the approximation error of the derivation procedure in (38). As is depicted in Fig. 6(c), the auto-correlation curve stands for the absolute value of $\overline{\bf R}_{\rm BS}[1,:]$, while the red circles show the value of the normalized sinc function. The approximation procedure shows negligible error, which confirms the high accuracy of our proposed approximation $(b)$ in (38).
@@ Appendix A Proof of Proposition 2
If the phase shift of reference wave is drawn from a uniform distribution, i.e., $\xi_{n}\sim{\mathcal{U}}[0,\pi]$, the interference term will vanish to 00 asymptotically, as $\mathbb{E}[\cos(2\xi_{n})]=0$ and $\mathbb{E}[\sin(2\xi_{n})]=0$ hold for $N_{\rm BS}\rightarrow\infty$.[37]

### v4

@@ II-C Sensing Model
${~{}~{}~{}}\!+\!\sum_{\ell=1}^{L}\!\sum_{m=1}^{N_{\rm UE}}\!\!\alpha_{\ell}{g}\!\left({\bf r}_{\rm S}^{(\ell)},{\bf r}_{\rm BS}^{(n)}\right)\!{g}\!\left({\bf r}_{\rm S}^{(\ell)},{\bf r}_{\rm UE}^{(m)}\right){\bf s}_{\rm S}[m]\!+\!{\bf e}_{\rm n}[n],$
where ${\bf s}_{\rm S}\in\mathbb{C}^{N_{\rm UE}}$ denotes the sensing signal, and ${\bf e}_{\rm n}\in\mathbb{C}^{N_{\rm BS}}$ is the additive white Gaussian noise (AWGN) vector. Without loss of generality, we consider the case when ${\bf s}_{\rm S}$ is an all-one vector [26]. As depicted in Fig. 3, the received object waves are then combined with the reference wave as ${\bf e}_{\rm ref}[n]+{\bf e}_{\rm obj}[n]$, where ${\bf e}_{\rm ref}=A[e^{\jmath\xi_{1}},\cdots,e^{\jmath\xi_{N_{\rm BS}}}]^{T}$ is the phase-shifted signal from the local oscillator with $A$ denoting the manipulated amplitude of the reference wave and $\xi_{n}$ denoting the phase shift at the $n$-th antenna element of the BS array. Then, the $n$-th power sensor [31, 32] deployed at the XL-MIMO transceiver measures the power pattern of combined signals as
@@ II-D Pilot Training Model
where ${\bf W}_{\rm RF}^{(t)}\in\mathbb{C}^{N_{\rm BS}\times N^{\rm RF}_{\rm BS}}$ and ${\bf W}_{\rm BB}^{(t)}\in\mathbb{C}^{N^{\rm RF}_{\rm BS}\times N^{\rm S}_{\rm BS}}$ denote the analog and digital combining matrices, and ${\bf F}_{\rm RF}^{(t)}\in\mathbb{C}^{N_{\rm UE}\times N^{\rm RF}_{\rm UE}}$ and ${\bf F}_{\rm BB}^{(t)}\in\mathbb{C}^{N^{\rm RF}_{\rm UE}\times N^{\rm S}_{\rm UE}}$ denote the analog and digital precoding matrices, respectively. Here, $N_{\rm UE}^{\rm RF}$ ($N_{\rm BS}^{\rm RF}$) and $N_{\rm UE}^{\rm S}$ ($N_{\rm BS}^{\rm S}$) denote the numbers of RF chains and data streams at the UE (BS), respectively, ${\bf n}^{(t)}\sim\mathcal{CN}(0,\sigma_{\rm n}^{2}{\bf I})$ is the AWGN vector, and $\sigma_{\rm n}^{2}$ is the power of the AWGN at each receive antenna.
@@ IV-A Time Inversion Near-Field Reconstruction
$+\left.\frac{e^{-2\jmath\xi_{n}-2\jmath\angle\overline{\alpha}_{\ell}+\jmath\kappa\left(\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|+\|{\bf r}-{\bf r}_{\rm BS}^{(n)}\|\right)}}{\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|}\right)$
$+\sum_{n=1}^{N_{\rm BS}}2{\rm Re}\left({\bf e}_{\rm n}[n]e^{\jmath\xi_{n}}\right)\frac{e^{\jmath\kappa\|{\bf r}-{\bf r}_{\rm BS}^{(n)}\|}}{\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|}$
@@
$\tilde{E}_{\ell}^{\rm d}({\bf r})$ $=\sum_{n=1}^{N_{\rm BS}}\frac{e^{-\jmath\kappa\left(\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|-\|{\bf r}-{\bf r}_{\rm BS}^{(n)}\|\right)}}{\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|},$ (18)
$\tilde{E}_{\ell}^{\rm i}({\bf r})$ $=\sum_{n=1}^{N_{\rm BS}}\frac{e^{-2\jmath\xi_{n}+\jmath\kappa\left(\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|+\|{\bf r}-{\bf r}_{\rm BS}^{(n)}\|\right)}}{\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|},$ (19)
@@
$\tilde{E}^{\rm n}({\bf r})=\sum_{n=1}^{N_{\rm BS}}2{\rm Re}\left({\bf e}_{\rm n}[n]\right)\frac{e^{-\jmath\xi_{n}+\jmath\kappa\|{\bf r}-{\bf r}_{\rm BS}^{(n)}\|}}{\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|}~{}~{}~{}~{}$ (20)
denote the desired term that contributes to the reconstruction, the interference term introduced by the power sampling procedure (7) of the $\ell$-th path, and the noise term, respectively.
@@
$\lim_{\begin{subarray}{c}\kappa\to{\infty}\\ N_{\rm BS}\to{\infty}\end{subarray}}\!\!\left|\tilde{E}_{\ell}^{\rm d}({\bf r})\right|\!=\!\begin{cases}\displaystyle\sum_{n=1}^{N_{\rm BS}}\frac{1}{d^{(\ell,n)}}\!\!\!\!&\text{if}~{}{\bf r}={\bf r}_{\rm S}^{(\ell)}\!,\\ 0&\text{otherwise},\end{cases}$ (23)
where $d^{(\ell,n)}=\|{\bf r}_{\rm S}^{(\ell)}-{\bf r}_{\rm BS}^{(n)}\|$, and the vanishing effect also holds for both the interference and noise terms as
@@
$\lim_{\begin{subarray}{c}\\ N_{\rm BS}\to{\infty}\end{subarray}}\left|\tilde{E}^{\rm n}({\bf r})\right|$ $=0,$ (25)
@@ IV-B Low-Complexity Time Inversion Algorithm
The locations of the UE and scatterers can be reconstructed exploiting the proposed time inversion method by enumerating possible target locations ${\bf r}$ in the near-field grid. Specifically, during the sensing stage, the BS array generates a uniformly distributed vector $\bm{\xi}=[\xi_{1},\cdots,\xi_{N_{\rm BS}}]$ and performs phase shift $e^{\jmath\xi_{n}}$ at the $n$-th antenna element. The near-field region is split into $G_{x}\times G_{y}$ grids, where $G_{x}$ ($G_{x}>N_{\rm BS}$) and $G_{y}$ denote the numbers of grids on the $x$-axis and the $y$-axis, respectively. The compensated wave at the $i$-th grid on $x$-axis and the $j$-th grid on $y$-axis, i.e., ${\bf r}^{(i,j)}=(x_{i},y_{j})$, is finally manipulated by the time inversion method using (14), and is then stored into the $i$-th row and the $j$-th column of the amplitude matrix $\bm{\Theta}\in\mathbb{C}^{G_{x}\times G_{y}}$, where
@@
which can be reformulated as the linear convolution of vector ¯$\overline{{\bm{\hbar}}}={{\bm{\hbar}}}\odot{\bf e}_{\rm ref}\in\mathbb{C}^{N_{\rm BS}}$ defined in (12) and $\overline{\bf g}_{j}=[e^{\jmath\kappa\sqrt{x_{1}^{2}+y_{j}^{2}}},\cdots,e^{\jmath\kappa\sqrt{x_{G_{x}}^{2}+y_{j}^{2}}}]^{T}\in\mathbb{C}^{G_{x}\times 1}$. Hence, for a fixed $j$-th grid on $y$-axis, the time inversion reconstructed result [^6] Convolution extends the dimension to $N_{\rm BS}+G_{x}-1$, which leads to incomplete reconstruction at the edges, and the valid results are obtained only for indices from $N_{\rm BS}$ to $G_{x}$. To satisfy the need for $G_{x}$ valid points, the $x$-axis grid will be extended to $G_{x}^{\prime}=G_{x}+2N_{\rm BS}$ to compensate the associated edge effects. can be recast as
@@
where $\overline{\bf g}_{j}=[e^{\jmath\kappa\sqrt{x_{1}^{2}+y_{j}^{2}}},\cdots,e^{\jmath\kappa\sqrt{x_{G_{x}^{\prime}}^{2}+y_{j}^{2}}}]^{T}\in\mathbb{C}^{G_{x}^{\prime}\times 1}$ with $G_{x}^{\prime}=G_{x}+2N_{\rm BS}$. Notably, (30) can be further implemented efficiently using FFT as
@@ IV-C Proposed Multipath Localization
$\hat{\alpha}_{\ell}=\overline{\alpha}_{\ell}\sum_{n=1}^{N_{\rm BS}}\frac{1}{d^{(\ell,n)}}.$ (32)
@@ VI-C Sensing-Enhanced CE
We firstly investigate the sparsification performance of the proposed dictionary. The channel sparse representations of the DFT dictionary, spherical wave dictionary, and proposed dictionary are compared in Fig. 6(a), where the sparse representation is obtained by $\tilde{\bf h}=\bm{\Psi}^{\dagger}{\bf h}$. As can be observed, the conventional DFT dictionary shows a severe energy leakage effect in the near-field region [13], which can be improved by the spherical wave dictionary sampled in the polar-domain. Meanwhile, the proposed DPSS-based eigen-dictionary entails the sparsest pattern among the three dictionaries. Note that the proposed dictionary is compensated by the matrix ${\bf D}_{{\rm BS}({\rm UE})}$ in (39) and therefore, the sparse representation contains no specific angular information. In addition, different from the DFT and spherical dictionaries, the non-zero support set appears in the first several indices since the SVD always sorts the non-zero singular values first. We further validate the approximation error of the derivation procedure in (40). As is depicted in Fig. 6(b), the auto-correlation curve stands for the absolute value of $\overline{\bf R}_{\rm BS}[1,:]$, while the red circles show the value of the normalized sinc function. The approximation procedure shows negligible error, which confirms the high accuracy of our proposed approximation $(b)$ in (40).
@@ Appendix A Proof of Proposition 2
If the phase shift of reference wave is drawn from a uniform distribution, i.e., $\xi_{n}\sim{\mathcal{U}}[0,\pi]$, the interference term will vanish to 00 asymptotically, as $\mathbb{E}[\cos(2\xi_{n})]=0$ and $\mathbb{E}[\sin(2\xi_{n})]=0$ hold for $N_{\rm BS}\rightarrow\infty$[37].
Because $\mathbf{e}_{\mathrm{n}}$ is an AWGN vector, ${\rm Re}\left(\mathbf{e}_{\mathrm{n}}[n]\right)$ is a Gaussian random variable with zero mean. When $N_{\mathrm{BS}}\to\infty$, the noise term $\tilde{E}_{\ell}^{\rm n}({\bf r})\to 0$ according to the law of large numbers. ∎